OpenStudy (anonymous):

what is quadratic function with only the real two zeros given? x=-1 and x=-3

OpenStudy (anonymous):

the reqd eq is (x+1)(x+3)=0 or x^2+x+3x+3=0 or x^2+4x+3=0

OpenStudy (anonymous):

thats not one of my answer choices

OpenStudy (anonymous):

any one else have an idea

OpenStudy (anonymous):

@ktnguyen1 ????

OpenStudy (anonymous):

I had the same answer with @matricked

OpenStudy (anonymous):

But my choices are y=x^4-4x^3-4x^2-4x-3 y=-x^4+4x^3+4x^2+4x+3 y=x^4+4x^3+3x^2+4x-4 y=x^4+4x^3+4x^2+4x+3

OpenStudy (anonymous):

is one of those equal to your answer

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

u asked for a quadratic or apolynomial of 4 degree

OpenStudy (anonymous):

u can use factor and reaminder theorem to determine ur answer..

OpenStudy (anonymous):

I dont know im just giving you my answer choices?

OpenStudy (anonymous):

Im confused and dont know which to pick

OpenStudy (anonymous):

or divide each of ur expression on the right(only in x) and then after obtaining another quadratic determine the nature of the result to pick ur reqd choice

OpenStudy (anonymous):

do you know what it is because i cant figure it out

OpenStudy (anonymous):

i mean to say divide each of ur answer (expression in x) by x^2+4x+3 and check the resultant quadratic to obtain the real or imaginary roots

OpenStudy (anonymous):

I have to try synthetic devision to find the answer the last choice is the answer

OpenStudy (anonymous):

I get it! Thanks sooo much!

OpenStudy (anonymous):

you r wlcm! good night im so sleepy!!!!!!

OpenStudy (anonymous):

good Night!!!

OpenStudy (anonymous):

Me too!!